# Trojan Model Detection Using Activation Optimization

### v2

@@ Abstract
Training machine learning models can be very expensive or even unaffordable. This may be, for example, due to data limitations (unavailability or being too large), or computational power limitations. Therefore, it is a common practice to rely on open-source pre-trained models whenever possible. However, this practice is alarming from a security perspective. Pre-trained models can be infected with Trojan attacks, in which the attacker embeds a trigger in the model such that the model’s behavior can be controlled by the attacker when the trigger is present in the input. In this paper, we present a novel method for detecting Trojan models. Our method creates a signature for a model based on activation optimization. A classifier is then trained to detect a Trojan model given its signature. We call our method TRIGS for TRojan Identification from Gradient-based Signatures. TRIGS achieves state-of-the-art performance on two public datasets of convolutional models. Additionally, we introduce a new challenging dataset of ImageNet models based on the vision transformer architecture. TRIGS delivers the best performance on the new dataset, surpassing the baseline methods by a large margin. Our experiments also show that TRIGS requires only a small amount of clean samples to achieve good performance, and works reasonably well even if the defender does not have prior knowledge about the attacker’s model architecture. Our dataset will be released soon.
@@ 2.3 Defenses against Trojan attacks
DeepInspect [3] is an algorithm for detecting models with backdoors assuming that the defender has access only to the trained model and no access to clean data samples. To achieve this goal, the method uses model inversion [9] to construct a training dataset for the model. Using the constructed training data, a generator is employed to create perturbation patterns (triggers) such that the model produces a given target class with the poisoned samples. Then, anomaly detection is applied to determine if any of the generated triggers is a real trigger used to train the model. Similarly, in [36][41], anomaly detection is applied to detect the real trigger (if any) among a set of generated triggers. The idea was extended in [6] to the black-box case, where the model is only accessible through its query responses. However, in this case, the triggers are generated by reverse engineering with real clean samples. In [21][31], potentially compromised neurons are first identified. Based on the identified compromised neurons, possible triggers are generated and only those that consistently subvert the model’s predictions to a certain target class are admitted. The methods require at least one clean sample of each class. The case when the clean samples available to the defender are limited or non-existent was handled in [38]. The method uses the similarity between two embeddings per image, one with a universal perturbation pattern and one with a local perturbation pattern, as an indicator of the presence of a Trojan. Universal Litmus Patterns (ULPs) [17] are input patterns to a classification model, the output of which can be used to distinguish benign from Trojan models. Few patterns per class have been shown to be effective in detecting Trojan models on multiple datasets. A very similar idea was proposed and thoroughly analyzed in [44]. Complex attack scenarios, in which the trigger pattern is not limited to be patch-shaped, were the focus of [22]. Our proposed method, TRIGs, is a white-box defense mechanism that can work well with a limited access to clean data. Unlike recent defenses, which are customized for a specific architecture [5], TRIGS is generic and works well with both CNN and ViT architectures.
@@ 3.1 Threat model
A similar threat model in the black box setting was used in [17, 44]. However, targeting the white box case is still important due to the wide-spread use of pretrained model weights downloaded from the web. In such cases, when the model weights are available, it is imperative to leverage them to enhance the detectability of Trojan models.
@@ 4.1 Evaluation data
| CIFAR10 | VGG | Train |
500 |
500 |
| CIFAR10 | VGG | Test |
100 |
100 |
| Tiny ImageNet | ResNet | Train |
1000 |
1000 |
| Tiny ImageNet | ResNet | Test |
100 |
100 |
| ImageNet | ViT | Train |
500 |
500 |
| ImageNet | ViT | Test |
100 |
100 |
@@ 4.3 Evaluation results
The two baseline methods, which are ULP [17] and k-Arm optimization [31], were chosen based on the availability of their code and its adaptability to new datasets. For ULP, we used the publicly available code for CIFAR10 and Tiny ImageNet. We applied the best configuration in the paper for each dataset, which was ten litmus patterns for CIFAR10 and five for Tiny ImageNet. We created ten sets of litmus patterns for each dataset. In Tab. 2, we report the mean and standard deviation of the AUC score over the litmus pattern sets. It is worth noting that we could not reproduce or even come close to the results reported in the ULP paper despite using the code released by the authors. For the ImageNet dataset, we adapted the public implementation for Tiny ImageNet to work with the ImageNet dataset. We experimented with different learning rates and numbers of litmus patterns. Generally, it was hard to get the model to converge. We obtained the best results with ten litmus patterns and the same training settings used for the Tiny ImageNet dataset in the authors’ code. Due to the heavy computational cost of creating litmus patterns for ImageNet, we only report the results of one set of litmus patterns.
@@
From the results in Tab. 2 and Figs. 6, 7 and 8, we can observe that, in each dataset, at least three of our four variants surpass the baseline performance, regardless of whether the probe model is CNN or ViT-based. Except for the activation maximization variant, all other variants improve over baseline methods by statistically significant large margins.
It is interesting to observe that the pixel-wise statistics variant is the best in the case of CIFAR10 and Tiny ImageNet. This variant achieves the best mean score and the lowest standard deviation. It is also comparable to the minimization variant in the case of ImageNet. However, for ImageNet, the variant that combines both types of activation optimization maps achieves the best performance. It is also interesting to notice that the activation minimization variant consistently performs better than the activation maximization one. This result is surprising given that all prior work on model inversion focused on activation maximization (or alternatively minimizing the classification loss, e.g. the cross-entropy loss). Here, for the first time, we find a good use for activation minimization-based model inversion.
| CIFAR10 | Tiny ImageNet | ImageNet | ||
| ULP | 0.64 (0.060) | 0.74 (0.075) | 0.58 | |
| k-Arm | 0.48 (0.008) | 0.56 (0.12) | 0.50 (0.67) | |
| TRIGS | Min | 0.96 (0.011) | 0.96 (0.015) | 0.82 (0.108) |
| Max | 0.60 (0.067) | 0.93 (0.016) | 0.73 (0.013) | |
| Both | 0.95 (0.022) | 0.98 (0.010) | 0.94 (0.015) | |
| Stats | 0.99 (0.003) | 1.00 (0.001) | 0.84 (0.046) |

### v3

@@ Abstract
Training machine learning models can be very expensive or even unaffordable. This may be, for example, due to data limitations (unavailability or being too large), or computational power limitations. Therefore, it is a common practice to rely on open-source pre-trained models whenever possible. However, this practice is alarming from a security perspective. Pre-trained models can be infected with Trojan attacks, in which the attacker embeds a trigger in the model such that the model’s behavior can be controlled by the attacker when the trigger is present in the input. In this paper, we present a novel method for detecting Trojan models. Our method creates a signature for a model based on activation optimization. A classifier is then trained to detect a Trojan model given its signature. We call our method TRIGS for TRojan Identification from Gradient-based Signatures. TRIGS achieves state-of-the-art performance on two public datasets of convolutional models. Additionally, we introduce a new challenging dataset of ImageNet models based on the vision transformer architecture. TRIGS delivers the best performance on the new dataset, surpassing the baseline methods by a large margin. Our experiments also show that TRIGS requires only a small amount of clean samples to achieve good performance, and works reasonably well even if the defender does not have prior knowledge about the attacker’s model architecture. Our code and data can be accessed through this page github.com/vimal-isi-edu/trigs.
@@ 2.3 Defenses against Trojan attacks
DeepInspect [3] is an algorithm for detecting models with backdoors assuming that the defender has access only to the trained model and no access to clean data samples. To achieve this goal, the method uses model inversion [9] to construct a training dataset for the model. Using the constructed training data, a generator is employed to create perturbation patterns (triggers) such that the model produces a given target class with the poisoned samples. Then, anomaly detection is applied to determine if any of the generated triggers is a real trigger used to train the model. Similarly, in [37][43], anomaly detection is applied to detect the real trigger (if any) among a set of generated triggers. The idea was extended in [6] to the black-box case, where the model is only accessible through its query responses. However, in this case, the triggers are generated by reverse engineering with real clean samples. In [22][32], potentially compromised neurons are first identified. Based on the identified compromised neurons, possible triggers are generated and only those that consistently subvert the model’s predictions to a certain target class are admitted. The methods require at least one clean sample of each class. The case when the clean samples available to the defender are limited or non-existent was handled in [40]. The method uses the similarity between two embeddings per image, one with a universal perturbation pattern and one with a local perturbation pattern, as an indicator of the presence of a Trojan. Universal Litmus Patterns (ULPs) [18] are input patterns to a classification model, the output of which can be used to distinguish benign from Trojan models. Few patterns per class have been shown to be effective in detecting Trojan models on multiple datasets. A very similar idea was proposed and thoroughly analyzed in [46]. Complex attack scenarios, in which the trigger pattern is not limited to be patch-shaped, were the focus of [23]. More recently [38], a detection method was introduced based on the observation that Trojan models have an anomalously large logit margin for the target class. Our proposed method, TRIGs, works both in the white-box and black-box settings and with a limited access to clean data. Unlike recent defenses, which are customized for a specific architecture [5], TRIGS is generic and works well with both CNN and ViT architectures. The closest defense to TRIGS is the One-Pixel Signature (OPS) defense [14], in which a model signature is used to train a binary classifier to distinguish Trojan from benign models. However, to work in the black-box scenario, OPS uses brute force search to construct the signatures instead of using gradient descent optimization as in TRIGS. Also, the signature size in OPS is proportional to the number of classes, which can be very large, while TRIGS can leverage pixel statistics to significantly reduce the signature size regardless of the number of classes.
@@ 3.1 Threat model
A similar threat model in the black box setting was used in [18, 14, 46]. We show that our approach still works in the black-box setting. However, it is important to note that targeting the white box case is still practical due to the wide-spread use of pretrained model weights downloaded from the web. In such cases, when the model weights are available, it is imperative to leverage them to enhance the detectability of Trojan models.
@@ 4.1 Evaluation data
CIFAR10 | VGG | Train | 500 | 500 |
CIFAR10 | VGG | Test | 100 | 100 |
Tiny ImageNet | ResNet | Train | 1000 | 1000 |
Tiny ImageNet | ResNet | Test | 100 | 100 |
ImageNet | ViT | Train | 500 | 500 |
ImageNet | ViT | Test | 100 | 100 |
@@ 4.3 Evaluation results
The three baseline methods, which are ULP [18], k-Arm optimization [32], and MM-BD [38], were chosen based on the availability of their code and its adaptability to new datasets. For ULP, we used the publicly available code for CIFAR10 and Tiny ImageNet. We applied the best configuration in the paper for each dataset, which was ten litmus patterns for CIFAR10 and five for Tiny ImageNet. We created ten sets of litmus patterns for each dataset. In Tab. 2, we report the mean and standard deviation of the AUC and accuracy scores over the litmus pattern sets. It is worth noting that we could not reproduce or come close to the results reported in the ULP paper despite using the code released by the authors. For the ImageNet dataset, we could not get the method to work due to excessive computational cost and lack of convergence.
@@
For the MM-BD method, we adapted the publicly available code to work with our datasets. We found that the default number of steps used in the paper (300) was too small for the models to converge. For a fair comparison, in all our experiments, we let the optimization run until convergence. We used ten different runs for each model in the CIFAR10 and the Tiny ImageNet datasets. However, due to the excessive computational time, we only used one run for the ImageNet dataset.
From the results in Tab. 2 and Figs. 6, 7 and 8, we can observe that, in each dataset, at least one of our four variants surpasses or matches the baseline performance, regardless of whether the probe model is CNN or ViT-based. Moreover, when our method surpasses the baseline methods, the margin is statistically significant.
It is interesting to observe that the pixel-wise statistics variant is the only variant that consistently outperforms or matches all baseline methods. It is also the best in the case of CIFAR10 and Tiny ImageNet, achieving the highest mean score and the lowest standard deviation. However, for ImageNet, the variant that combines both types of activation optimization maps achieves the best performance. It is also interesting to notice that the activation minimization variant consistently performs better than the activation maximization one. This result is surprising given that all prior work on model inversion focused on activation maximization (or alternatively minimizing the classification loss, e.g. the cross-entropy loss). Here, for the first time, we find a good use for activation minimization-based model inversion.
Out of the three baseline methods, the only serious contender is the MM-BD method. In fact, this method achieves a perfect AUC score on the Tiny ImageNet dataset (though its accuracy is not the best). However, similar to the other two baseline methods, MM-BD struggles on the ImageNet dataset. We believe this struggle is due to the ViT architecture, in which the main assumption of the MM-BD method (the presence of an anomalously large logit margin for the target class in a Trojan model) may not hold.
CIFAR10 | Tiny ImageNet | ImageNet | |||||
AUC | Acc | AUC | Acc | AUC | Acc | ||
ULP | 0.64 (0.060) | 0.61 (0.048) | 0.74 (0.075) | 0.71 (0.066) | – | – | |
k-Arm | 0.68 (0.028) | 0.51 (0.000) | 0.65 (0.120) | 0.54 (0.024) | 0.51 (0.67) | 0.5 (0.000) | |
MM-BD | 0.90 (0.012) | 0.79 (0.029) | 1.00 (0.000) | 0.97 (0.009) | 0.59 | 0.51 | |
TRIGS | Both | 0.95 (0.022) | 0.90 (0.038) | 0.98 (0.010) | 0.93 (0.014) | 0.94 (0.015) | 0.87 (0.033) |
Max | 0.60 (0.067) | 0.57 (0.054) | 0.93 (0.016) | 0.83 (0.047) | 0.73 (0.013) | 0.66 (0.020) | |
Min | 0.96 (0.011) | 0.92 (0.019) | 0.96 (0.015) | 0.92 (0.013) | 0.82 (0.108) | 0.75 (0.083) | |
Stats | 0.99 (0.003) | 0.96 (0.011) | 1.00 (0.001) | 0.99 (0.010) | 0.84 (0.046) | 0.76 (0.050) |